# GenCO: Generating Diverse Solutions to Design Problems with Combinatorial Nature

### v1

@@ 4 Experiments
We tested our approach on two applications: level generation for the Zelda game and map generation for path planning in Warcraft. Although both applications are in the game domain, they have quite different use cases. Nevertheless, both settings involve combinatorial optimization in the pipeline, which makes the typical use of deep generative models nontrivial. In these experiments, we demonstrated that GenCO significantly outperforms the baselines, efficiently finding diverse, realistic solutions that obey combinatorial properties. This success paves the way for combining combinatorial optimization with deep generative models.
4.1 Generating game levels using GANs
We evaluate our model on the task of generating diverse Zelda game levels. In this setting, we use GenCO to train using a GAN-like constrained formulation (refer to section 3.2 and Appendix B).
4.1.1 Settings
Explicitly Constrained GAN: Game Level Design
@@ 4.1.1 Settings
We train GenCO on the task of generating Zelda game levels. Here, we are given examples of human-crafted game levels and are asked to generate fun new levels. The generated levels must be playable in that the player must be able to solve them by moving the character through a route that reaches the destination. Additionally, the levels should be realistic in that they should be somewhat similar to the real game levels as measured by an adversary that is trained to distinguish between real and fake images ($\mathcal{L}_{\text{gen}}{}$). We use the same dataset as (Zhang et al., 2020), consisting of 50 Zelda game levels, as well as the same network architectures, as we don’t tune the hyperparameters for our model in particular but rather compare the different approaches, all else equal.
@@
Here, we evaluate several approaches, including the previous work (Zhang et al., 2020), which we call GAN + MILP fix. This approach first trains a standard Wasserstein GAN architecture to generate the game levels. Specifically, they train the WGAN by alternatively training two components: a generator and a discriminator. The generator is updated to “fool” the discriminator as much as possible in that it tries to maximize the loss of the discriminator. The discriminator tries to correctly separate the generated and real game levels into their respective classes. When the practitioner then wants to generate a valid level, this approach generates a level with the generator and then fixes it using a MILP formulation that finds the nearest feasible game level where proximity is determined by cosine distance, which is equivalent to minimizing a dot product distance.
@@
We evaluate two variants of GenCO, GenCO - Fixed Adversary, which iteratively updates the generative model to fool a fixed pretrained adversary, and GenCO - Updated Adversary, which updates both the generator and adversary during training. The fixed adversary approach trains to fool a pretrained adversary that is obtained from the fully trained GAN from previous work (Zhang et al., 2020). Both GenCO approaches are initialized with the fully trained GAN from previous work (Zhang et al., 2020).
@@ 4.1.2 Results
Approach % unique
GAN loss ($\mathcal{L}_{\text{gen}}{}$)
(lower better)
GenCO adversary
(lower better)
GenCO - Updated Adversary 0.995 -10.10 -4.49
GenCO - Fixed Adversary 0.22 -1.45 -0.85
GAN + MILP fix (Previous Work) 0.52 0.22 0.24
Table 1: Game level design comparison. We compare GenCO with an updating adversary against GenCO with a fixed adversary, and lastly, previous work that postprocesses solutions to be valid game levels. Here, all levels are valid, but GenCO exhibits more diversity and realism.
GAN + MILP Fix Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GenCO Fixed Adv Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GenCO Updated Adv Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Generated zelda game examples. The GenCO Updated Adversary obtains solutions that seem more realistic than mainly empty fixed GAN instances and fixed adversary instances.
We present a table of results in Table 1 and examples in Figure 2, which includes the performance of the previous approach as well as two variants of our proposed approach. In these settings, we estimate performance based on sampling 1000 levels. Each level is made out of a grid, with each grid cell having one of 8 components: wall, empty, key, exit, 3 enemy types, and the player. A valid level is one that can be solved by the player in that there is a valid route starting at the player’s location, collecting the key, and then reaching the exit. We evaluate the performance of the models using two types of metrics: diversity, as measured by the percentage of unique levels generated, and fidelity, as measured by the average objective quality of a fixed GAN adversary. We evaluate using adversaries from both the previous work and GenCO. As shown in Table 1, we find that GenCO with an updating adversary generates unique solutions at a much higher rate than previous approaches and also generates solutions that are of higher quality as measured by both the GAN adversary and its own adversary. The adversary quality demonstrates that the solutions are realistic in that neither the adversary from the previous work nor from GenCO is able to distinguish the generated examples from the real examples. This is further demonstrated in Figure 2 with the updated adversary generating realistic and nontrivial game levels. Furthermore, given that the levels are trained on only 50 examples, we can obtain many more game levels. Note here that both approaches are guaranteed to give playable levels as they are postprocessed to be valid. However, GenCO is able to generate more diverse solutions that are also of higher quality.
@@
As shown in Table 1, GenCO, with an updated adversary, obtains the highest percentage of unique solutions, generating 995 unique solutions out of 1000. This is significantly higher than the previous work, which only generated 520 unique solutions. This is likely due to the fact that the generator is trained with the downstream fixing explicitly in the loop. This means that while the previous work may have been able to “hide” from the adversary by generating slightly different continuous solutions, these continuous solutions may project to the same discrete solution. On the other hand, by integrating the fixing into the training loop, GenCO’s generator is unable to hide in the continuous solution space and thus is heavily penalized by generating the same solution as the adversary will easily detect those to be originating from the generator. In essence, this makes the adversary’s task easier as it only needs to consider distinguishing between valid discrete levels rather than continuous and unconstrained levels. This is also reflected in the adversary quality, where GenCO’s adversary is able to distinguish between levels coming from the previous work’s generator and the real levels with a much better loss.
@@ 4.2 Map generation for path planning
In this experiment, we consider somewhat different and more challenging for GenCO setting of generation of image maps for strategy games like Warcraft. This poses a greater challenge because the generator $G$ described in equation 1 is not restricted and can produce any image of a map (setting descrbied in section 3.3). However, we are required to encourage $G$ to generate a map with two main criteria: 1) it should resemble real and diverse game maps, as determined by an adversary trained to differentiate between real and fake images; 2) the cost of the shortest path from the top-left to the bottom-right corners (source and destination) should be minimized. Intuitively, it corresponds to (mostly) a diagonal part of the image, and a map can include various elements (terrains) like mountains, lakes, forests, and land, each with a specific cost (for example, mountains may have a cost of 3, while land has a cost of 0). To calculate this, we pass the generated image to the fixed ResNet (called “cost NN”) to get the graph representation together with edge costs. The cost NN corresponds to the mapping $g(\cdot)$ in Algorithm 2. This is then followed by the shortest path solver ($h(\cdot)$ in Algorithm 2). The objective is to populate the map with a diverse range of objects while ensuring the shortest path remains low-cost. We use the same dataset as in (Pogančić et al., 2020) with DCGAN architecture adapted from (Zhang et al., 2020). More implementation details and experimental settings can be found in Appendix A.
@@ 4.2.1 Baselines
We examine various baseline models, including an “Ordinary GAN” that does not incorporate the Shortest Path objective (see the top part of Fig. 1). Like the Zelda experiment, this approach employs a standard Wasserstein GAN architecture closely following (Zhang et al., 2020). However, in this case, we generate images directly instead of encoding game levels. More precisely, we train the WGAN by iteratively training two components: a generator and a discriminator. The generator is fine-tuned to deceive the discriminator to the greatest extent possible, aiming to maximize the discriminator’s loss. The discriminator, on the other hand, endeavors to accurately distinguish between generated and authentic game levels and categorize them accordingly.
The next baseline is “GAN + cost NN”, which simultaneously feeds the generator’s output into both the discriminator and a ResNet that calculates the costs for each edge in the grid. We then average the output to obtain the final loss for generator. We then average these outputs to derive the ultimate loss for the generator. While this approach considers costs associated with objects, it does not incorporate information about the Shortest Path.
@@ 4.2.2 Results
Quantitative results are showcased in Table 2. The “Ordinary GAN” focuses exclusively on the GAN’s objective, without considering the Shortest Path’s objective $f$. In contrast, GenCO achieves higher performance with regards to the SP’s objective, albeit with a slight reduction in GAN’s loss. On the other hand, the “GAN + cost NN” approach optimizes using the cost vector of the entire map (generated by ResNet). In other words, it tries to uniformly avoid placing costly objects in any part of the image, whereas we are interested only in the shortest path. While it demonstrates a modest enhancement in the objective $f$, it experiences a notable decline in terms of generator loss ($\mathcal{L}_{\text{gen}}$). This indicates a trade-off between optimizing for the Shortest Path and the GAN’s objective.
Approach
GAN loss ($\mathcal{L}_{\text{gen}}$)
(lower better)
SP loss ($\mathcal{C}$)
(lower better)
Ordinary GAN 0.6147 36.45
GAN + cost NN 0.8994 35.61
GenCO (ours) 0.6360 23.99
Table 2: Performance comparison on Warcraft map generation. Our here goal is to create a map that is both realistic (“fools” discriminator, i.e. lower GAN loss) and has small objective for the shortest path (SP). Results are averaged across 100 instances.
@@
Such quantitative results directly translate into image qualities. In fig. 3, a subset of generated Warcraft map images using various methods is displayed. The “Ordinary GAN” tends to generate maps with elements such as mountains and lakes, which are considered “very costly”, particularly along the Shortest Path from the top-left to the bottom-right (which mostly goes through the diagonal). This makes sense since WGAN is trained with no information about grid costs. In contrast, the “GAN + cost NN” approach produces less costly maps, albeit with reduced diversity. For instance, most part of the image is populated with the same object. On the other hand, GenCO strikes a balance, achieving a cheap Shortest Path while maintaining a diverse range of elements on the map. The Shortest Path is efficient, and the map exhibits a rich variety of features simultaneously.
Ordinary GAN Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GAN + cost NN Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GenCO (ours) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 3: A subset of generated Warcraft map images using various methods. “Ordinary GAN” generates “very costly” maps (e.g. mountains, lakes) along the Shortest Path from top-left to bottom-right; “GAN + cost NN” generates less costly maps but they are less diverse; GenCO: SP path is cheap and the map is diverse at the same time.

### v2

@@ 4 Experiments
We test GenCO on three applications: Zelda game level design, Warcraft path planning map generation, and diverse inverse photonic device design. Settings are summarized in Table 2, and all settings involve combinatorial optimization, making the application of deep generative models nontrivial. In these experiments, GenCO significantly outperforms the baselines, efficiently finding diverse, realistic solutions that obey combinatorial properties, paving the way for combining combinatorial optimization with deep generative models.
4.1 Metrics
We evaluate generative performance with density and coverage (Naeem et al., 2020). Density measures, for the average fake data point, how many real examples are nearby. Coverage measures how many real data points have a fake data point nearby. Formally, $M$ fake examples $Y_{j}$, are evaluated against $N$ real examples $X_{i}$. The metrics use the ball around $X_{i}$, with radius $r=\text{NND}_{k}(X_{i})$ being the maximimum distance from $X_{i}$ to its $k$ nearest real neighbors, to quantify generated density and coverage of the real data distribution. They are expressed as density = $\frac{1}{kM}\sum_{j=1}^{M}\sum_{i=1}^{N}1_{Y_{j}\in B(X_{i},\text{NND}_{k}(X_{i}))}$, and coverage = $\frac{1}{N}\sum_{i=1}^{N}1_{\exists j\text{s.t.}Y_{j}\in B(X_{i},\text{NND}_{k}(X_{i}))}$.
4.2 Explicitly Constrained GAN: Game Level Design
Approach % Unique $\uparrow$ Density $\uparrow$ Coverage $\uparrow$ GAN loss ($\mathcal{L}$) $\downarrow$ GenCO adversary $\downarrow$
GAN + MILP fix (previous) 0.52 0.07 0.94 0.22 0.24
GenCO - Fixed Adversary 0.22 0.05 0.98 -1.45 -0.85
GenCO - Updated Adversary 0.995 0.06 0.82 -10.10 -4.49
Table 3: Game level design comparison between the previous MILP postprocessing work, GenCO with an updating adversary, and GenCO with a fixed adversary, with all approaches guaranteed to generate feasible levels. We measure % Unique Density and Coverage to estimate diversity and distribution alignment. We also evaluate the adversary’s prediction to gauge how well a neural network can distinguish the generated levels from real levels.
GAN+MILP Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GenCO fixed Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GenCO upd Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Generated zelda level examples. “GenCO Updated” obtains solutions that seem more realistic than the empty MILP postprocessed GAN levels and GenCO Fixed Adversary levels.
@@ 4.2 Explicitly Constrained GAN: Game Level Design
We evaluate GenCO on generating diverse and realistic Zelda game levels. Automatic level generation with playability guarantees helps designers create games with automatically generated content or iterate on development using a few examples. In this setting, we use GenCO to train using a constrained WGAN formulation. Here, we are given examples of human-crafted game levels $\mathcal{X}_{\mathrm{train}}$ and are asked to generate fun new levels. The generated levels must be playable in that the player must be able to complete them by moving the character through a route that reaches the destination $({\bm{x}}\in\Omega)$. Additionally, the levels should be realistic in that they should be similar to the real game levels as measured by an adversary that is trained to distinguish between real and fake images. This adversary imposes the generative group loss $\mathcal{L}{}$, effectively approximating the Wasserstein distance (Arjovsky et al., 2017). We use the same 50 Zelda levels as (Zhang et al., 2020), and the same neural network architecture, as we don’t tune the hyperparameters for our model in particular but rather compare the different approaches, all else equal.
4.2.1 Baselines
@@
Here, we evaluate GenCO against previous work (Zhang et al., 2020)(GAN + MILP fix). This approach trains a standard Wasserstein GAN to approximate human game levels. Specifically, they train the WGAN by alternatively training a generator and a discriminator. The generator is updated to “fool” the discriminator (maximize the loss of the discriminator). The discriminator tries to correctly separate the generated and real game levels into their respective classes. When the practitioner wants to generate a valid level, this approach generates a soft level and fixes it using a MILP that finds the nearest feasible game level in terms of cosine distance (dot product).
@@
We evaluate two variants of GenCO, GenCO - Fixed Adversary, which trains against a fixed pretrained adversary from previous work (Zhang et al., 2020), and GenCO - Updated Adversary, which updates both the generator and adversary during training. Both approaches are initialized with the fully trained GAN from previous work (Zhang et al., 2020). Importantly, the adversaries act on the generated solution ${\bm{x}}\in\Omega$, rather than the latent soft solution ${\bm{c}}$.
@@ 4.2.2 Results
Results are presented in Table 3 and examples in Figure 2, which includes the performance of the previous approach as well as two variants of GenCO. In these settings, we estimate performance based on sampling 1000 levels. Each level is made out of a grid, with each grid cell having one of 8 components: wall, empty, key, exit, 3 enemy types, and the player. A valid level is one that can be solved by the player in that there is a valid route starting at the player’s location, collecting the key, and then reaching the exit. We evaluate the performance of the models using two types of metrics: diversity, as measured by the percentage of unique levels generated, and fidelity, as measured by the average objective quality of a fixed GAN adversary. We evaluate using adversaries from both the previous work and GenCO.
As shown in Table 3, we find that GenCO with an updating adversary generates unique solutions at a much higher rate than previous approaches and also generates solutions that are of higher quality as measured by both the GAN adversary and its own adversary. The adversary quality demonstrates that the solutions are realistic in that neither the adversary from the previous work nor from GenCO is able to distinguish the generated examples from the real examples. This is further demonstrated in Figure 2 with the updated adversary generating realistic and nontrivial game levels. Furthermore, given that the levels are trained on only 50 examples, we can obtain many more game levels. Overall, GenCO - Updated Adversary obtains the best uniqueness and generative loss. However, it obtains neither the best density nor coverage since game levels are generally sparse (the majority of tiles are empty floors), meaning that having a slightly shifted map will result in a higher distance than a completely empty map. As a result, coverage and density have counterintuitive behaviors in discrete settings while in standard image domains, samples are continuous and unconstrained. Additionally, density and coverage are relatively low. Note here that both approaches are guaranteed to give playable levels as they are postprocessed to be valid. However, GenCO is able to generate more diverse solutions that are also of higher realism quality.
Approach Density $\uparrow$ Coverage $\uparrow$ GAN loss ($\mathcal{L}$) $\downarrow$ SP loss ($\mathcal{D}$) $\downarrow$
Ordinary GAN 0.81 0.98 0.6147 36.45
GAN + semantic loss (Di Liello et al., 2020) 1.09 0.98 0.8994 35.61
GenCO (ours) 0.94 0.93 0.6360 23.99
Table 4: Map generation for Warcraft performance comparison on 100 instances. The goal is to generate maps which are similar to the data distribution (Density / Coverage), realistic (lower GAN loss or group loss $\mathcal{L}$), and have small shortest paths (SP or individual loss $\mathcal{D}$).
@@
As shown in Table 3, GenCO, with an updated adversary, obtains the highest percentage of unique solutions, generating 995 unique solutions out of 1000, significantly higher than the 520 unique solutions in previous work. This is likely due to the fact that the generator is trained with the downstream fixing explicitly in the loop. This means that while the previous work may have been able to “hide” from the adversary by generating slightly different continuous solutions, these continuous solutions may project to the same discrete solution. On the other hand, by integrating the fixing into the training loop, GenCO’s generator is unable to hide in the continuous solution space and thus is heavily penalized by generating the same solution as the adversary will easily detect those to be originating from the generator. In essence, this makes the adversary’s task easier as it only needs to consider distinguishing between valid discrete levels rather than continuous and unconstrained levels. This is also reflected in the adversary quality, where GenCO’s adversary is able to distinguish between levels coming from the previous work’s generator and the real levels with a much better loss.
@@ 4.3 Map generation for path planning
In this experiment, we consider somewhat different and more challenging for GenCO setting of generation of image maps for strategy games like Warcraft. This poses a greater challenge because the generator $G$ is not restricted and can produce any RGB image of a map (group loss $\mathcal{L}$ operates directly on latent space $\mathcal{C}$, i.e., $\mathcal{L}(\mathcal{C})$ in eq. 3). However, we would like to encourage $G$ to generate a map with two main criteria: 1) it should resemble real and diverse game maps, as determined by the group Loss $\mathcal{L}(\mathcal{C})$; 2) the cost of the shortest path (SP) from the top-left to the bottom-right corners (source and destination) should be minimized (group loss $\mathcal{D}({\bm{g}}({\bm{c}}))$). Intuitively, the SP corresponds to (mostly) a diagonal part of the image, and a map can include various elements (terrains) like mountains, lakes, forests, and land, each with a specific cost (for example, mountains may have a cost of 3, while land has a cost of 0). To calculate this, we pass the generated image to the fixed ResNet to get the graph representation together with edge costs. This is then followed by the shortest path solver (${\bm{g}}$). The objective is to populate the map with a diverse range of objects while ensuring the shortest path remains low-cost: $\mathcal{L}(\mathcal{C})+\gamma\sum_{j}\mathcal{D}({\bm{g}}({\bm{c}}_{j}))$. We use the same dataset as in (Pogančić et al., 2020) with DCGAN architecture adapted from (Zhang et al., 2020). Note that the maps are unconstrained, representing the generator’s latent space. However, the generated feasible solution is a valid path through the map. Additionally, the generative loss acts on the latent space $c$ rather than the feasible solution as we are given maps rather than solutions. The resnet architecture is part of the problem formulation in that it determines costs. More implementation details and experimental settings can be found in Appendix A.
@@ 4.3.1 Baselines
We examine various baseline models, including an “Ordinary GAN” that does not incorporate the Shortest Path objective. As in the Zelda experiment, this approach employs a standard Wasserstein GAN architecture closely following (Zhang et al., 2020). However, in this case, we generate images directly instead of encoding game levels. More precisely, we train the WGAN by iteratively training two components: a generator and a discriminator. The generator is fine-tuned to deceive the discriminator to the greatest extent possible, aiming to maximize the discriminator’s loss. The discriminator, on the other hand, endeavors to accurately distinguish between generated and authentic game levels and categorize them accordingly.
The next baseline is a “soft” penalty approach: “GAN + semantic loss”. It uses a fixed NN to extract representation of the image in the form of cost vector (obtained from the fixed ResNet). This vector is then averaged and added to the final objective as a regularization. The concept is similar to the semantic loss appeared in (Di Liello et al., 2020). The difference is that instead of building a circuit (via knowledge compilation) to encode the constraints, we use a fixed pre-trained ResNet as a surrogate to obtain the cost vector. While this baseline considers costs associated with objects, it does not incorporate information about the Shortest Path.
@@ 4.3.2 Results
Quantitative results are showcased in the Table 4. The “Ordinary GAN” focuses exclusively on the GAN’s objective (group loss $\mathcal{L}$), without considering the Shortest Path’s objective (individual loss $\mathcal{D}$). In contrast, GenCO achieves higher performance with regards to the SP’s objective, albeit with a slight reduction in GAN’s loss. On the other hand, the “GAN + semantic loss” approach tries to uniformly avoid placing costly objects in any part of the image, whereas we are interested only in the shortest path. While it demonstrates a modest enhancement in the objective $\mathcal{D}$, it experiences a notable decline in terms of group loss ($\mathcal{L}$). This indicates a trade-off between optimizing for the Shortest Path and the GAN’s objective. Here, the density and coverage are comparable between all approaches. While GenCO doesn’t obtain the highest density and coverage, it does obtain good SP loss and comparable GAN loss with Ordinary GAN, indicating a trade-off between fitting the distribution and obtaining high-quality solutions.
@@
Such quantitative results directly translate into image qualities. In fig. 3, a subset of generated Warcraft map images using various methods is displayed. The “Ordinary GAN” tends to generate maps with elements such as mountains and lakes, which are considered “very costly”, particularly along the Shortest Path from the top-left to the bottom-right (which mostly goes through the diagonal). This makes sense since WGAN is trained with no information about grid costs. In contrast, the “GAN + semantic loss” approach produces less costly maps, albeit with reduced diversity. For instance, most part of the image is populated with the same object. On the other hand, GenCO strikes a balance, achieving a cheap Shortest Path while maintaining a diverse range of elements on the map. The Shortest Path is efficient, and the map exhibits a rich variety of features simultaneously.
$\gamma$ SP Loss ($\mathcal{D}$) $\downarrow$ Density $\uparrow$ Coverage $\uparrow$
0 36.45 0.81 0.98
1e-4 27.66 0.93 0.93
1e-3 23.99 0.94 0.93
3e-3 23.76 0.75 0.86
5e-3 18.02 0.49 0.61
1e-2 0.00 0.00 0.00
Table 5: $\gamma$ ablation, trading off individual and generative loss. High trade-off parameters improve optimization performance while low trade-off values give slightly better generation.
Additionally, in Table 5 we evaluate sensitivity to generative and shortest path loss trade-off parameter. While the shortest path loss improves with high trade-off parameter, the image quality metrics start to deteriorate rapidly. Eventually, our approach generates blank or empty images, as such images result in the shortest path with zero cost. This behavior could be attributed to the combinatorial nature of the problem.
Ord. GAN Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GAN + sem. loss Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GenCO Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 3: A subset of generated Warcraft map images. “Ordinary GAN” generates “very costly” maps (e.g. mountains, lakes) along the Shortest Path from top-left to bottom-right; “GAN + semantic loss” generates less costly maps but they are less diverse; GenCO: SP path is cheap and the map is diverse at the same time.
4.4 Constrained VAE: Photonic Design
Approach % Unique $\uparrow$ Density $\uparrow$ Coverage $\uparrow$ Avg Solution Loss $\downarrow$
VQVAE + postprocess 30.6% 0.009 0.006 1.244
GenCO (reconstruction only) 100% 0.148 0.693 1.155
GenCO (objective only) 46.6% 0.013 0.036 0
GenCO (reconstruction + objective) 100% 0.153 0.738 0
Table 6: Inverse photonic design comparison evaluating variants of GenCO using VQVAE against the same model architecture which postprocesses solutions. Solution loss is evaluated using a simulation of Maxwell’s equations and is optimal at 0.
We finally evaluate GenCO for training a Vector Quantized Variational Autoencoder (VQVAE) to ensure that all decoded examples satisfy combinatorial manufacturing constraints ${\bm{x}}\in\Omega$, are each minimizers of a photonic loss $\mathcal{D}$, and are similar to given solutions via the ELBO loss $\mathcal{L}$. We evaluate GenCO against a generative + postprocess baseline (Zhang et al., 2020) on the inverse photonic design setting.
The inverse photonic design problem (Schubert et al., 2022) asks how to design a device consisting of fixed and void space to route wavelengths of light from an incoming location to desired output locations at high intensity. Here the feasible region consists of satisfying manufacturing constraints that a die with a specific shape must be able to fit in every fixed and void space. Specifically, the fixed and void regions must represent the union of the die shape at different locations. The objective function here consists of a nonlinear but differentiable simulation of the light using Maxwell’s equations. Previous work demonstrated an approach for finding a single optimal solution to the problem. However, we propose generating a diverse collection of high-quality solutions using a dataset of known solutions. In practice, there are various applications for generating multiple feasible photonic devices. Firstly, there are many devices that globally minimize the objective function by matching the specification perfectly. With a generator, practitioners can select from a diverse array of solutions that already optimize both specification and satisfy manufacturing constraints, quickly generating more examples if needed. Upon inspecting the solutions, they may select one that satisfies their needs, or derive high-level insights from the diversity of viable solutions to improve the design process. Lastly, optimizing the encoded objective function over the feasible region may not necessarily result in the best solution considering unquantified components like aesthetics, further manual scrutiny, and simply tacit domain expertise. Regardless, the selected design must always be optimal and satisfy the problem constraints. As such, we seek to generate diverse optimal solutions to aid the practitioner when a single solution fails to capture the full range of possibilities.
In this setting, we train a vector quantized variational autoencoder (VQVAE) (Van Den Oord et al., 2017). During training, the encoder encodes a known solution and then decodes an object with the same shape as the input space, using neural networks for the encoder and decoder. The continuous decoded object is then fed into the constrained optimization layer to enforce that the generated solution is feasible. This feasible solution is then used to calculate the reconstruction loss. Furthermore, in this setting we have an optimization-based objective function that consists of the simulation of Maxwell’s equations. As such, we ablate the data distribution approximation, and penalty components of GenCO: whether or not to train using the reconstruction loss, and whether or not to penalize generated solutions based on Maxwell’s equations. We consider a baseline of training the same generative architecture without combinatorial optimization and then postprocessing generated examples during evaluation using a combinatorial solver. At test time, we sample noise and decode to the original input space. We then obtain a feasible solution using the combinatorial solver. The results in Table 6 include the solution uniqueness rate, the average loss evaluated using Maxwell’s equations, as well as the density and coverage. The dataset is obtained by expensively running previous work (Ferber et al., 2023) until we obtain 100 optimal solutions. We evaluate performance on generating 1000 feasible examples.
These results in Table 6 demonstrate that postprocessing obtains very few unique solutions which all have high loss and furthermore don’t resemble the data distribution. This is largely due to the method not being trained with the postprocessing end-to-end. Although it closely approximates the data distribution with continuous and infeasible objects, when these continuous objects are postprocessed to be feasible, they are no longer representative of the data distribution and many continuous solutions collapse to the same discrete solution. Additionally, GenCO obtains high density and coverage compared to the baseline postprocessed VQVAE indicating that training end-to-end helps in this setting.
GenCO - reconstruction only gives many unique solutions that resemble the data distribution. However, the generated devices fail to perform optimally in the photonic task at hand. Disregarding the reconstruction loss and only training the decoder to generate high-quality solutions yields high-quality but redundant solutions. Combining both the generative reconstruction penalty as well as the nonlinear objective, GenCO generates a variety of unique solutions that optimally solve the inverse photonic design problem and are representative of the data distribution.